News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 16 November 2018 9.00am EST

During the last week Chandra completed the observing schedule as planned.

Real-time procedures were executed on Nov 10 to dump OBC-A and -B memory as a follow-up to the patches uplinked last week. The dumps will be used to update the baseline memory images maintained on the ground.

A Chandra image release was issued on Nov 15 describing the combination of Chandra X-ray observations and LOFAR low-frequency radio observations of the cluster of galaxies Abell 1033. The data show that the cluster is composed of two subclusters in the process of merging, and reveal a variety of highly energetic phenomena, including shock waves in the hot X-ray emitting intracluster gas produced from the high speed of the merging process, and jets emerging from several of the cluster galaxies. The image also shows a "radio phoenix" consisting of a cloud of electrons that faded in radio emission but was then reenergized when shock waves compressed the cloud. For details see: http://chandra.si.edu/photo/2018/a1033/

The schedule of targets for the next week is shown below and includes an observation of ESO 338-IG04, which was accepted as a Director's Discretionary Time Target of Opportunity on Sep 21, and an observation of MAXIJ1820+070, which was accepted as a Target of Opportunity on Nov 14 and is coordinated with the VLA.



------------------------------------------
        Radiation Belts
        ESO338-IG04HLX-1       ACIS-S       Nov 19
        MAXIJ1820+070          ACIS-S/HETG
        SDSSJ172227.18+3207    ACIS-S
        Perseus_west_cf_2      ACIS-I
        PKS0405-123            HRC-S /LETG  Nov 20
        PSRJ0656-2228          ACIS-S
        Abell2219              ACIS-I
        Radiation Belts                     Nov 21
        Perseus_west_cf_2      ACIS-I
        PKS0405-123            HRC-S /LETG  Nov 22
        Abell2219              ACIS-I       Nov 23
        Radiation Belts                     Nov 24
        Abell2219              ACIS-I
        SN2012fh               ACIS-S
        RWAur                  ACIS-S
        LSBCF570-06            ACIS-S       Nov 25
        PSRJ1838+1650          ACIS-S
        A1795                  ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports